OpenStudy (anonymous):

Prove the the following equation has a value of 2.

OpenStudy (anonymous):

Prove that**

OpenStudy (anonymous):

\[(4+\sqrt{15})(\sqrt{10}-\sqrt{6})\sqrt{4-\sqrt{15}}\]

OpenStudy (anonymous):

let me use my graph tablet :)

OpenStudy (anonymous):

Hah, alright

OpenStudy (anonymous):